WLCG-OSG-EGEE Ops' Minutes Mon 06 Oct 2008

Summary

CREAM CE has been released to production, but will not yet be selected by default WMS match-making. Next week, many gLite 3.0 services will be made obsolete.

Attendance

The meeting started at 16:04. We should really try to start at 16:00 out of respect to those who make the effort to connect on time.

The list of attendees is taken directly from web conference display. Please ensure that you open a browser window and enter your name and affiliation - even if you also dial in by phone!

EGEE & WLCG

  • OCC / CERN ROC: John Shade, Antonio Retico, Nick Thackray, Steve Traylen, Diana Bosio
  • Harry Renshall (WLCG Service Coordination)
  • Miguel Coelho Dos Santos (CERN Site)
  • Lev Shamardin (Russian ROC)
  • Ron Trompert
  • Victor Edneral, Russian COD team
  • Jeremy Coles (UKI)
  • Torsten Antoni (GGUS-KIT)
  • Kostas Koumantaros SEE
  • Malgorzata Krakowian (CE ROC)
  • Gert Svensson NE ROC
  • Matt Hodges (RAL-LCG2)
  • Kai Neuffer (SWE/PIC)
  • 0033478930880

OSG Representatives

  • Tim Silvers (Indiana)
  • John McGee RENCI (present at CERN)
  • Catalin Dimitrescu (FNAL)

LHC Experiments

  • ATLAS: Alessandro di Girolamo
  • LHCb: Roberto Santinelli
  • CMS: absent
  • ALICE: absent

Feedback on Last Week's Minutes

None was given.

EGEE Items

Grid Operator Hand Over on Duty

  Primary Team Secondary Team
From ROC France ROC Italy
To ROC UKI ROC Russia

Issues mainly from AP region, some UK, and one at CERN.

  • Does ROC UKI receive notifications from GGUS? Jeremy will follow-up (wasn't aware of any problems).
  • Russia will follow-up the APEL problem on one of their sites
  • Steve will personally follow-up with AP ROC

PPS Reports

  • CREAM CE pilot is entering a new phase. The plan is to setup 20 CEs, 3 Italian sites and 1 German site currently involved, others would be welcome.
  • PPS received two releases this week (mainly removal of obsolete packages and FTS bug-fixes)
  • update 33 is being prepared with the ability to publish tool versions

gLite Release News

  • gLite 3.1 update 32 released to production. The big news is CREAM CE (an older version than the one used in PPS pilot), which is the recommended version for ALICE. It will not be matched by default WMS.
  • gLite 3.0 update 34 released with a new VOMS server host certificate
  • for more details, follow the links on the agenda page

EGEE Items From ROC Reports

  • ROC France asked about progress on GGUS:40565. John answered that the incorrect reporting of MAINT status was caused by timing problems due to the gap between when tests get submitted and when they publish their results. SAM is working on fixing this by saving the maintenance status when tests are submitted in the results DB. Tests will stop returning MAINT as a status, but all visualisation tools will be able to display a consistent picture! We're hoping to have this fix implemented in the next fortnight.
  • GSTAT is now comparing LFC services comparing Information System and GOCDB. The tests are insignificant in terms of availability and COD alarms; the aim is solely to increase accuracy.

New LFC SAM tests

  • Explained in the agenda. Roberto said that LHCb needed special LFC tests. John replied that if they were generic, the SAM team would provide them, but since they were specific to LHCb, LHCb should develop their own tests.

gLite 3.0 services to be obsoleted

  • Check the list! No new updates will be released, and you'll be on your own. Please check, because most gLite 3.0 services are in the list, and it's happening next week.

WLCG Items

  • On the issue of VOs making direct requests to sites to install software, Steve underlined the policy that things must first be discussed in the Ops meeting, and then mentioned in the VO Card. ATLAS agrees to this procedure (they need libblas library). A two-week timeline will be set, and if a site is unhappy, they can raise it in the following week's meeting.
  • Steve apologized for the jobs that he ran for GridFest, as he should have sent a broadcast first. A CE in Lancaster broke, several hundred thousand jobs ran, and the RTM demo was very successful.

WLCG issues coming from ROC reports

  • On the subject of France wondering whether there was a procedure to be notified of changes in the LHC DN alarm list, Torsten said that sites don't need to know, but they could consult the Wiki if they so desired. Since the mails are signed by GGUS, the justification for the request had disappeared.

Upcoming WLCG Service Interventions

  • Please check the URL in the agenda page for the many interventions this week.

ATLAS Service

  • No report submitted, nothing special to report

ALICE Service

  • No one present

CMS Service

  • No one present

LHCb Service

Roberto repeated the report that he'd given to the 3 o'clock meeting.
  • Problem at CERN getting enough slots to run (fair share to be investigated).
  • For the LHCb service master at CERN and Tier1s, only Gridka has broadcast information about tomorrow's 1hr downtime.
  • Need to review the rules for generating gridmap file. All users in pool should be mapped to a single local user account. Steve said that gridmap file is not supported everywhere and is to become a thing of the past. Roberto will write things down and send to John [his input is now attached to original meeting agenda in LHCb report section].

Storage Service

  • New Castor versions and SRM interfaces (of interest to RAL and Taiwan).

OSG Items

  • Tim has been testing new ticketing exchange system. Things are going well, and manual steps should be eliminated. This seemed to please Maria, who nevertheless brought up three tickets which were discussed:
    • GGUS:37059 - has been fixed according to John McGee, and should be closed
    • GGUS:39303 - no activity, so Maria suggests to close
    • GGUS:40225 - also looks fairly obsolete
  • Alessandro wondered whether OSG handles Tier2 issues within the cloud differently that in Europe (by going directly to BNL, the Tier1). Maria replied that this was entirely up to them.

AOB

  • Alessandro had a follow-up query from Wednesday's GDB as to whether the direct assignment of Team tickets to Tier2 would be taken out. "Not at all", replied Steve.
  • Catalin from Fermilab complained that there had been no follow-up to problems with the BDII and "no compatible resources". Steve requested a written reminder of what the issue was.
  • Maria advertised the VOMS Admin/VOMRS workshop (21st-24th October) and invited all interested parties to attend. It's F2F, but a phone connection can be arranged.

Meeting ended at 17:00

Action Items

Newly Created Action Items

Assigned to Due date Description State Closed Notify  
Main.OCC 2007-03-05 Example Action Item 2007-03-06 SteveTraylen   edit

Review of Open Action Items

Open Action Items

IdSubmitterDescriptionCreationDueAssigned To 

Actions Closed in Last 20 Days

IdSubmitterDescriptionCreationDueAssigned ToClosed 

Next Meeting

The next meeting will be Monday, 13 OCT 2008 15:00 UTC (16:00 Swiss Mountain Time).

  • Attendees can join from 14:45 UTC (15:45 Swiss local time) onwards.
  • The meeting will start promptly at 15:00 UTC (16:00 Swiss local time).
  • The WLCG section will start at the fixed time of 15:30 UTC (16:30 Swiss local time).
  • To dial in to the conference:
    • Dial +41227676000
    • Enter access code 0157610


These minutes can only be changed by members of:

Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2008-10-10 - JohnShade
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    EGEE All webs login

This site is powered by the TWiki collaboration platform Powered by Perl This site is powered by the TWiki collaboration platformCopyright & by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Ask a support question or Send feedback